A blood pressure test was given to 450 women ages 20 ton36. It showed that their mean systilic blood pressure was 118.0 mm HG, with a standard deviation of 12.9 mm Hg.
(A) Determine the z-score, to the nearest hundredth, for woman who had a systilic blood pressure trading of 114.1 mm Hg.
(B) The z-score for one woman was 2.18. What was her systilic blood pressure trading? (Round your answer to one decimal place.)

Answer by Boreal(15235)

z=(x-mean)/sd
z=(114.1-118)/12..9
z=-0.30
b. 2.18=(x-118)/12.9
28.12=x-118
x=146.1 mm systolic